A study shows that length of the careers of professional basketball players are nearly normally distributed, with a mean of 6.1 years and a standard deviation of 1.8 years. If there are 250 players, how many have a career of less than 5 years?

1
Expert's answer
2022-01-18T10:33:03-0500

"z=\frac{x-\mu}{\sigma}=\frac{5-6.1}{1.8}=-0.61"


"P(z<-0.61)=0.2709"


number of students who have a career of less than 5 years:

"n=0.2709\cdot250=67.725\approx 68" students